# OneForecast: A Universal Framework for Global and Regional Weather Forecasting

### v1

@@ E.3 Typhoon Tracking
This study focuses on two extreme cyclones: Tropical Storm Yagi and Severe Typhoon Molave. The Yagi formed near Ito To, Japan on August 6, 2018, and landed over Wenling, China on August 12. The Molave formed on October 11, 2020, and landed over the Philippines on October 25, 2020. The initial conditions for these two cyclones are set at 0:00 UTC, August 6, 2018 and 0:00 UTC, 11 October 2020, respectively. Since there is no Fuxi results in 2018 on weatherbench2, we can not compare it with Yagi.
Appendix F Related Work
F.1 Deep Learning based Weather Forecasting
F.2 Numerical analysis methods
@@ Appendix G Additional Results
We present more additional results in Figure 8, 9, 10 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, including 18 variables that are importmant to weather forecasting, each with results ranging from 6 hours to 10 days. These additional results further demonstrate the effectiveness of OneForecast. Same as the Figure 3 , the initial conditions is 00:00 UTC, 1 January 2020.

### v2

@@ F.3 Typhoon Tracking
This study focuses on two extreme cyclones: Tropical Storm Yagi and Severe Typhoon Molave. The Yagi formed near Ito To, Japan on August 6, 2018, and landed over Wenling, China on August 12. The Molave formed on October 11, 2020, and landed over the Philippines on October 25, 2020. The initial conditions for these two cyclones are set at 0:00 UTC, August 6, 2018 and 0:00 UTC, 11 October 2020, respectively. Since there is no Fuxi results in 2018 on WeatherBench2, we can not compare it with Yagi. The results of Best Track (Ying et al., 2014) (Lu et al., 2021) can be found in https://tcdata.typhoon.org.cn/en/.
@@ Appendix G Additional Results
G.1 Efficiency Analysis
As shown in Table 5, our OneForecast has a competitive performance for Parameters and MACs. For the MACs, the size of input tensor is set to (1, 69, 120, 240). Not that for the ML-based weather forecasts, the computational cost is less important compared with the forecasting accuracy because the ML-basd model is several orders of magnitude faster (maybe tens of thousands of times) than traditional numerical methods. For instance, in numerical forecasting, a single simulation for a 10-day forecasts can take hours of computation in a supercomputer that has hundreds of nodes. In contrast, ML-based weather forecasting models just need a few seconds or minutes to produce 10-day forecasts using only 1 GPU.
G.2 Spectral Analysis
As shown in Figure 8, we compute the surface kinetic energy spectrum and Q700 spectrum for baseline models using WeatherBench2’s official results (averaged across the first 700 ICs). Our OneForecast model achieves comparable performance in this standardized evaluation framework. Notably, as Q700 data for Fuxi were not available in the WeatherBench2, only its surface kinetic energy spectrum could be analyzed.
Figure 8: Spectral analysis of different models.
G.3 Quantitative analysis of Extreme Event
To assess the forecast performance of more extreme events, as shown in Table 6, we present 2 extreme event assessment indicators (the higer the better) CSI and SEDI. And we also add a quantitative metric for typhoon prediction in Table 7, a lower value represents better results. It can be seen that our OneForecast also achieves satisfactory results in quantitative analysis.
Table 6: Quantitative analysis of extreme event for different models.
      Model       Wind10M       Wind10M       T2M       T2M
      CSI       SEDI       CSI       SEDI
      Pangu       0.11       0.29       0.16       0.34
      Graphcast       0.13       0.29       0.20       0.38
      Fuxi       0.11       0.20       0.19       0.27
      Ours       0.14       0.31       0.21       0.40
Table 7: Quantitative analysis of typhoon for different models.
      Model       Track Position Error(km)
      FS-HRES       332
      Pangu 1.5°       222
      Graphcast 1.5°       212
      Pangu       231
      Graphcast       197
      Ours       157
Figure 9: Additional quantitative comparison with the results from WeatherBench2 for several important variables.
G.4 Additional quantitative comparison with the results from WeatherBench2
As shown in Figure 9, we compare all models released by WeatherBench2, except for ENS (ensemble forecasting, not the same task) and Spherical CNN (too few ICs, only 178 compared with our used 700). While the WeatherBench2 baseline leverages numerous training strategies, we only conducted 1-epoch of finetuning (Ours_finetune) during the brief rebuttal period. Nevertheless, a 2-epoch finetune model (Ours_finetune2) demonstrates improved results, indicating the potential for further gains with additional finetuning. If we finetuned for more epochs, OneForecast can achieve better result. However, our primary objective is to introduce a novel paradigm for global and regional weather forecasting rather than solely optimizing metrics, we just finetune for a few epoch as an example.
G.5 Additional Visual Results
We present more additional results in Figure 10, 11, 12 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, including 18 variables that are importmant to weather forecasting, each with results ranging from 6 hours to 10 days. These additional results further demonstrate the effectiveness of OneForecast. Same as the Figure 3, the initial conditions is 00:00 UTC, 1 January 2020.